To find : No of protons in 10gm of calcium carbonate

 To find :
 
No of protons in 10gm of calcium carbonate

Grade:11

1 Answers

Arun
5 years ago
Dear Ashok
 
3*10^24 protons 
as 100 gm CaCO3 =6.022*10^23 molecules of CaCO3, 
so 1 gm CaCO3 = 6.022*10^21 molecules, 
also 1 molecule of CaCO3 contains 50 protons i.e,20+6+24 
so no. of protons in 10 gm of CaCO3 is =6.022*10^21*10*50 =3.011*10^24 protons.
 
Regards
Arun